# Intents, Techniques, and Components: a Unified Analysis of Interaction Authoring Tasks in Data Visualization

### v2

@@ 9.3. Generative Power
First of all, both Interaction+ and Lyra 2 focus on creating interactivity for marks, and mostly overlook the other types of visual objects such as collection and scene. By including these visual objects in the scope, techniques such as OnSet, scrollytelling and dynamic toggling are possible. Second, it is important that an expressive interaction authoring language or tool needs to understand and provide internal abstractions for various representation and interaction components. It would also be helpful to expose these components in the interface through manipulable visual representations by adopting, for example, a visual programming paradigm (Kuhail et al., 2021). Finally, a powerful authoring tool can allow users to start at any of the three levels of abstraction, depending on their expertise and tasks, and intelligently provide assistance through appropriate abstractions. The introduction mentioned one example where an intent can be translated into techniques and low-level implementations. We can imagine a different scenario where an author has composed an interaction technique using certain components, and wants to see how the implementation can be generalized to other techniques and even intents.
@@ Acknowledgements.
This research was carried out as part of the requirements for Hyemi Song’s Master’s degree in Computer Science at the University of Maryland, College Park from 2022 Spring to 2023 Summer.

### v3

@@ 9.3. Generative Power
First of all, both Interaction+ and Lyra 2 focus on creating interactivity for marks, and mostly overlook the other types of visual objects such as collection and scene. By including these visual objects in the scope, techniques such as OnSet, scrollytelling and dynamic toggling are possible. Second, it is important that an expressive interaction authoring language or tool needs to understand and provide internal abstractions for various representation and interaction components. It would also be helpful to expose these components in the interface through manipulable visual representations by adopting, for example, a visual programming paradigm (kuhail_characterizing_2021). Finally, a powerful authoring tool can allow users to start at any of the three levels of abstraction, depending on their expertise and tasks, and intelligently provide assistance through appropriate abstractions. The introduction mentioned one example where an intent can be translated into techniques and low-level implementations. We can imagine a different scenario where an author has composed an interaction technique using certain components, and wants to see how the implementation can be generalized to other techniques and even intents.
@@ Acknowledgements.
This research was conducted as part of the requirements for Hyemi Song’s Master’s degree in Computer Science at the University of Maryland, College Park, from Spring 2022 to Summer 2023. This paper was peer-reviewed and invited for fast-track submission to the Computer Graphics Forum (CGF) journal via EuroVis 2023.
Contact
For correspondence, please contact the first author (hsong02@cs.umd.edu) at the University of Maryland College Park.